1. Electron and proton are placed at two different locations (one charge at each location). The locations have the same electric potential. Do these charges have the same electric potential energy?

a. Yes, if the electric potentials at the two locations are the same, the electric potential energies are also the same, regardless of the sign and magnitude of the charges placed at these locations.

b. No, because the electric potential energy EPE at a given location depends on the charge placed at that location as well as the electric potential V.

c. Both A and B

d. None of the above

2. Two electric charges are separated by 1.5 m (q1=-3.70Β΅C and q2=+5.10Β΅C). Find the electric potential midway between them.

a. 9.11π‘₯104𝑉

c. 7.11π‘₯104𝑉

b. 8.11 π‘₯104𝑉

d. 6.11π‘₯104𝑉



1
Expert's answer
2022-03-06T15:18:45-0500

1. b.


2. The electric potential midway between them is given as follows:


"V = k\dfrac{q_1}{r} + k\dfrac{q_2}{r} = k\dfrac{q_1+q_2}{r}"

where "r = 1.5m\/2 = 0.75m" is the distance from each charge to the midway point. Thus, obtain:


"V = 9\times 10^9\cdot \dfrac{(-3.7 + 5.10)\times 10^{-6}}{0.75} \approx 1.7\times 10^4V"

Answer. "1.7\times 10^4V".